A person drives to a destination at a rate of thirty mph and returns over the same route at fifty mph. How far is the destination if the time returning is one hour less than the time going?
Answer by Alan3354(69443)

A person drives to a destination at a rate of thirty mph and returns over the same route at fifty mph. How far is the destination if the time returning is one hour less than the time going?
===============
d = r*t
d = 50t = 30(t+1)
50t = 30t + 30
t = 1.5 hours
d = 50*1.5 = 75 miles